OpenStudy (lena772):

Choose the equation below that represents the line passing through the point (2, -4) with a slope of one half. y = one halfx + 5 y = one halfx - 3 y = one halfx - 5 y = one halfx + 3

ganeshie8 (ganeshie8):

point-slope form

ganeshie8 (ganeshie8):

IF, Point : \((x_1, y_1)\) Slope : \(m\) then, line : \(y-y_1 = m(x-x_1)\)

OpenStudy (lena772):

y--4=m(x-2)

OpenStudy (lena772):

y+4=m(x-2)

ganeshie8 (ganeshie8):

yes, plugin \(m\) also

OpenStudy (lena772):

There are no x2 and y2 values to solve m.

ganeshie8 (ganeshie8):

we dont need to solve \(m\) \(m\) is given to us, read ur question again :o

ganeshie8 (ganeshie8):

... with a slope of one half. slope = 1/2 \(m\) = 1/2

OpenStudy (lena772):

Oh, sorry. y+4=1/2(x-2)

ganeshie8 (ganeshie8):

yes, expand right side.. . and change it to the form y = mx + b

OpenStudy (lena772):

y+4=1/2x-1

OpenStudy (lena772):

y=1/2x-5

ganeshie8 (ganeshie8):

\(\large \color{red}{\checkmark}\)